# Survey on Publicly Available Sinhala Natural Language Processing Tools and Research

### v25

@@ 3.11 Sentiment Analysis
A simple MLP-based method to classify sentiments in Sinhala text was initially proposed by Medagoda [273] based on their prior work [274]. A word2vec based tool118118118http://bit.ly/2QKI9Np for sentiment analysis of Sinhala news comments is available. A methodology for constructing a sentiment lexicon for Sinhala Language in a semi-automated manner based on a given corpus was proposed by Chathuranga et al. [275]Demotte et al. [276] proposed a sentiment analysis system119119119https://github.com/theisuru/sentiment-tagger based on sentence-state LSTM Networks for Sinhala news comments. In the subsequent work [277], they used word similarity to generate a Sinhala semantic lexicon. They followed this up with a further study [246] which discussed a number of other deep learning models such as RNN and Bi-LSTM in the domain of Sinhala sentiment analysis. Jayasuriya et al. [278] proposed a method to classify Sinhala posts in the domain of sports into positive and negative class sentiments. Ranathunga and Liyanage [279] claimed that using word embedding models as semantic features can compensate for the lack of well developed language-specific linguistic or language resources in the case of analysing sentiment of Sinhala news comments. Jayasuriya et al. [280] conducted a comparative study between word N-grams and character N-grams in the task of semantic classification of Sinhala content in social media. Which they soon followed up with an ensemble approach [281]. The work by Karunarathne [282] used word embedding to analyse the sentiment of manually annotated Sinhala Tweets. Abeyratne and Jayaratne [283] conducted a multi model analysis on classifying Sinhala songs by emotion.
@@
The work by Jayawickrama et al. [284, 285] used the data set released by Wijeratne and de Silva [99] to predict the reactions induced by Sinhala Facebook posts. They then extended the work [286] and compared the results obtained with their data set against that of Senevirathne et al. [246]Aththanayaka and Naleer [287] used Random forest, Support vector machines, and Multinomial Naïve Bayes models to analyse sentiment in Sinhala-English code-mixed text from social media. The ACTSEA dataset120120120https://bit.ly/3oZoziI for Sinhala sentiment analysis was introduced by Jenarthanan et al. [135]. The dataset contains 318,308 Sinhala tweets annotated with emotions. The study by Dhananjaya et al. [288] uses a sentiment lexicon of a high‐resource language to fine-tune Pre‐trained multilingual language models (PMLMs) such as mBERT and XLM‐R on an intermediate task which in turn is then used on the sentiment classification task of Sinhala. A sentiment-annotated dataset of Sinhala-English code mixed YouTube comments has been created by Uthpala and Thirukumaran [289]. However, given that even their paper is behind a paywall, there is no free and public way to access this dataset. The work by Gurgurov et al. [230] compared the performance of LLMs for the Sinhala sentiment analysis task. They report the highest results for the Masked Language Modeling objective for Sinhala in both of their configurations where they used a mixture of ConceptNet (CN) and/or Wikipedia (wiki) language adapters in conjunction with Task Adapters (TA). Liyanaarachchi et al. [290] used character-level embedding to classify sentiments of Sinhala news comments as positive or negative. They claim that the 1-of-m embedding outperforms Log-m embedding for this task. Bandaranayake and Usoof [291] used BERT, DistilBERT, RoBERTa, and XLM-R to analyse the sentiment of Sinhala news comments [246]. They report that XLMR-large model achieves the best result. Fernando and Ranathunga [292] in their study on improving cross-lingual representation of multilingual language models for low-resource languages using linguistic entity masking used Sinhala code mixed sentiment analysis as one of the evaluations. The work by Peiris and Sumanathilaka [293] analyses emotions in Sinhala songs both in the directions of lyrics and audio. In their analysis of lyrics, they use RoBERTa to represent the Sinhala words. Rizvi et al. [294] used XLM-RoBERTa to extract keywords and classify aspects in Sinhala Code-Mixed data. Then in their follow up work [295] they again used XLM-RoBERTa, this time for sentiment analysis and, on top of that used SHAP [296] and LIME [240] to make the model explainable.
@@ 3.12 Hate Speech Detection
A machine learning approach to detect hate speech in Sinhala was proposed by De Silva [298]. A feature model and a data set121121121https://bit.ly/3FTNaMZ for Sinhala hate speech detection for youtube was proposed by De Saa and Ranathunga [128]Sandaruwan et al. [299] have attempted to identify abusive Sinhala comments in social media using text mining and machine learning techniques. A cyberbullying comment classification study for Sinhala was conducted by [300] where they used classical machine learning algorithms. The study by Hettiarachchi et al. [301] used classical machine learning methods to detect hate speech in Romanized Sinhala social media posts. While the basic idea is the same, they have avoided mentioning transliteration in their paper. The study by Samarasinghe et al. [302] proposed using CNNs for detecting hate speech in Sinhala text. Kariyawasam [303] proposed a machine learning approach for identifying toxic Sinhala language on social media. Guruge et al. [304] used an ensemble of Naïve Bayes, Support Vector Machine, XGBoost, MLP, and AdaBoost to detect hate speech in 49019 Tweets they collected from February of 2021 to April of 2021. Sandaruwan et al. [305] collected 3000 Sinhala comments corpus and conducted Multinomial Naïve Bayes to detect hate speech. Sheran [306] claims to have used machine learning to detect hate speech written in Sinhala or Singlish on social media. A deep learning approach for detecting hate speech in Sinhala tweets was explored by Munasinghe and Thayasivam [307]. The study by Shalinda and Munasinghe [308] utilized classical machine learning techniques such as Linear Support vector classifier, Random Forest Classification, SGD classifier, Logistic Regression, XGBoost classifier and multinomial Naive Bayes classifier on both Sinhala and Singlish (Romanized Sinhala) to identify hate words. They report that the SGD classifier over TF-IDF with uni-grams and bi-grams gives them the highest accuracy. Gamage et al. [309] conducted a comparative analysis on a number of embedding systems as well as classical frequency-based methods for Sinhala hate speech detection. The study by Fernando et al. [310] also claims to use machine learning and deep learning to detect hate speech in Sinhala. The study by Perera et al. [311] predicted Sinhala hate speech using user behaviour on twitter by applying ensemble Machine Learning to classify them by the five big personality traits. The follow-up study by Perera et al. [129] analyses Sinhala hate speech propagation on Twitter. They have also published a dataset122122122https://github.com/Isurie/Text-Classification-Module/tree/master/Dataset of 1600+ annotated Sinhala tweets. Rajapaksha et al. [312] used deep learning to identify trending periods of hate topics on Twitter in Sinhala. Arachchi et al. [313] used a web-based tool to translate between Sinhala and English to detect inappropriate word usage in Sinhala. The study by Ehelepola [314] analyzed Sinhala text from social media and e-commerce sites for hate speech. The work by Dikwatta et al. [315] used supervised algorithms to detect Sinhala hate speech in text from image posts on the Internet (memes). Wickramaarachchi et al. [316] used LSTM on BART to compare the title and description of Sinhala youtube videos to their audio in order to determine if they contain hate speech. The study by Muthuthanthri and Smith [317] used an annotated dataset from Facebook with TF-IDF and fastText embedding features with SVM, CNN, RNN, LSTM as well as BERT and GPT2. They concluded that BERT obtains the best result for detecting hate speech in English-Sinhala code-mixed data. Chavinda and Thayasivam [172] have used Multilingual Large Language Models (MLLMs) with Dual Contrastive Learning (DCL) to detect Sinhala hate speech on Facebook123123123https://huggingface.co/datasets/krishan-CSE/Facebook_Sinhala_Hate_Speech and Twitter124124124https://huggingface.co/datasets/krishan-CSE/Twitter_Sinhala_Hate_Speech data sets. They report TwHIN-BERT-base [318] model to have obtained the best results. The work by Jahnavi and Chaturvedi [319] used BERT with CNN and LSTM to detect hate speech in Sinhala. Perera et al. [320] uses XLM-RoBERTa to classify Sinhala youtube video titles and comments to severity levels of hate. This, by definition is a multi-class classification. The work by Tanjim [321] analyses multiple low-resourced languages in the context of Cyberbullying detection. Ultimately, for the case of Sinhala as the target language, they recommend using Assamese, Bengali, Nepali, Urdu, Hindi, or Bodo as the source language for transfer learning this task.
, title=Multilingual Cyberbullying Detection in Low-Resource Languages and Dialects
@@
After the release of the SOLD dataset by Ranasinghe et al. [140] for the hate speech detection task in Sinhala, a number of subsequent works have been conducted using it. Fernando and Deng [322] used traditional machine learning algorithms to predict Sinhala hate speech. Ranasinghe and Zampieri [323] compared the results of pre-trained mT5 on Sinhala tweets from the SOLD dataset against results for German, Spanish, Hindi, Korean, and Marathi. The work by Bestgen [324] compared the result of classical machine learning algorithms on simple n-gram features for the hate speech detection task on the SOLD dataset for Sinhala against results for Assamese and Bengali. Dmonte et al. [325] used the SOLD data set to determine whether machine translation impacts the performance of hate speech detection. Haturusinghe et al. [326] introduces a series of models named Subasa that are tested on the SOLD dataset. Their trained PMLs125125125https://github.com/haturusinghe/subasa-plm and LLMs126126126https://github.com/haturusinghe/subasa-llm are separately publicly available. Ou6t of these models, they report that their Subasa-XLM-R model outperforms even the state-of-the-art commercial LLMs when evaluated on the SOLD data set. Ranasinghe et al. [327] used the SOLD dataset for the Hate Speech And Offensive Content (HASOC) identification task127127127https://hasocfire.github.io/hasoc/2023/task1.html they introduced. In completing the task, the study by Narayan et al. [328] compared the effectiveness of a number of pre-trained deep learning models for Sinhala hate speech detection. They showed that XLM Roberta Base achieves a macro-f1 of 83.49% over the LSTM baseline of 75.30%. For the same task, Ojo et al. [329] used mBERT to identify hate and offensive content in Sinhala. The work by Rostamkhani and Eetemadi [330] compared the zero-shot performance of ChatGPT on the SOLD dataset against a number of other Sinhala models. They conclude that Zero-shot ChatGPT perform better in Sinhala hate speech detection than models that were not fine-tuned on SOLD. Dmonte et al. [141] used the SOLD data set to verify the new MT-Offense128128128https://github.com/LanguageTechnologyLab/MT-Offense data set they created by using NMT to translate the English offensive language benchmark dataset, OLID [142]. They report the best F1 of 0.550 for the translation done using nllb-200-3.3B [134].
@@ 3.13 Fake News Detection
A dataset consisting of Sinhala documents drawn from Sri Lankan news websites was published by Jayawickrama et al. [126] along with the benchmark misinformation classification models. A hybrid approach to detect Sinhala fake news on Social media was proposed by Wijayarathna and Jayalal [331] where the text content of the post is checked against credible sources and the authenticity of the user account posting the relevant post is evaluated by a rule-based points allocation schema. Wijayarathna and Jayalal [332] collected a set of 120 fake news tweets and 250 non-fake news tweets which they then converted to vectors by taking the fasttext vector for words and averaging. The vector representations of the tweets were then compared to predict whether the news containing it is real or fake. The study by Udurawana et al. [333] proposes to use an accuracy score (obtained by analysing the news content) and a credibility score (obtained by using a scoring mechanism) to detect fake news in Sinhala text. They also incorporate a module that classifies on the basis of passive aggressiveness. The study by Adihetti and Jayalal [334] used autoencoders to detect Sinhala fake news from social media posts. The study by Wickramaarachchi et al. [316] detected fake content in Sinhala youtube videos by comparing their title and description to the audio using LSTM on BART.
@@ 3.14 Word Sense Disambiguation
There have been multiple attempts to do word sense disambiguation (WSD) [335, 336, 337, 338, 339] for Sinhala. For this, Arukgoda et al. [192] have proposed a system named Aruth based on the Lesk Algorithm[340]. An online tool129129129http://aruth.herokuapp.com/, an API130130130https://bit.ly/3sJEYbS of the algorithm, and code along with data on github131131131https://github.com/jseanm1/aruthSWSD are available. For the same task, Marasinghe et al. [341] have proposed a system based on probabilistic modeling. A dialogue act recognition system which utilizes simple classification algorithms has been proposed by Palihakkara et al. [342]. A word sense disambiguation tool named Sinsense was introduced by Subasingha [343]. They used cross-lingual sense disambiguation where English sense disambiguation was used to obtain Sinhala sense disambiguation. However, neither their tool nor their full research is publicly available.
@@ 3.15 Text Summarizing
A deterministic process flow for automatic Sinhala text summarizing was proposed by Welgama [344]. The study by Wimalasuriya [345], which has the same name as the above work by Welgama [344], uses the graph-based TextRank algorithm for automatic Sinhala text summarizing. The use case of Sinhala Text summarization for government gazettes was explored by Jayawardane [346]. The study by Rathnayake et al. [347] compared the results of extractive and abstractive summarization on Sinhala textbooks. The study by Jahan and Wijesekara [348] compared the abstractive summarizing methods of TF-IDF and Text-Rank for Sinhala using ROUGE as the evaluation score. They concluded TF-IDF to be the superior choice. Hasan et al. [151] introduced the XL-sum132132132https://github.com/csebuetnlp/xl-sum/tree/master dataset which includes 3,414 Sinhala documents and their summaries collected from Sinhala BBC website133133133https://www.bbc.com/sinhala. The M3LS134134134https://github.com/anubhav-jangra/M3LS multi-modal dataset introduced by Verma et al. [152] contains 10,148 Sinhala documents, relevant images, and their summaries collected from Sinhala BBC website135135135https://www.bbc.com/sinhala. Further, they claim that MT5 [349] obtains the best ROUGE-1 and ROUGE-L f scores for their Sinhala data set. The first Sinhala multi-document summarizing dataset, M2DS136136136https://huggingface.co/datasets/KushanH/m2ds, was created by Hewapathirana et al. [153]. The M2DS dataset consists of 23.5k Sinhala documents in 5.5k clusters with each cluster having a golden summary. Thier code and pre-trained models are also available on github137137137https://github.com/KushanMH/m2ds.

### v26

@@ 3.11 Sentiment Analysis
A simple MLP-based method to classify sentiments in Sinhala text was initially proposed by Medagoda [286] based on their prior work [287]. A word2vec based tool143143143http://bit.ly/2QKI9Np for sentiment analysis of Sinhala news comments is available. A methodology for constructing a sentiment lexicon for Sinhala Language in a semi-automated manner based on a given corpus was proposed by Chathuranga et al. [288]Demotte et al. [289] proposed a sentiment analysis system144144144https://github.com/theisuru/sentiment-tagger based on sentence-state LSTM Networks for Sinhala news comments. In the subsequent work [290], they used word similarity to generate a Sinhala semantic lexicon. They followed this up with a further study [257] which discussed a number of other deep learning models such as RNN and Bi-LSTM in the domain of Sinhala sentiment analysis. Jayasuriya et al. [291] proposed a method to classify Sinhala posts in the domain of sports into positive and negative class sentiments. Ranathunga and Liyanage [292] claimed that using word embedding models as semantic features can compensate for the lack of well developed language-specific linguistic or language resources in the case of analysing sentiment of Sinhala news comments. Jayasuriya et al. [293] conducted a comparative study between word N-grams and character N-grams in the task of semantic classification of Sinhala content in social media. Which they soon followed up with an ensemble approach [294]. The work by Karunarathne [295] used word embedding to analyse the sentiment of manually annotated Sinhala Tweets. Abeyratne and Jayaratne [296] conducted a multi model analysis on classifying Sinhala songs by emotion.
@@
The work by Jayawickrama et al. [297, 298] used the data set released by Wijeratne and de Silva [102] to predict the reactions induced by Sinhala Facebook posts. They then extended the work [299] and compared the results obtained with their data set against that of Senevirathne et al. [257]Aththanayaka and Naleer [300] used Random forest, Support vector machines, and Multinomial Naïve Bayes models to analyse sentiment in Sinhala-English code-mixed text from social media. The ACTSEA dataset145145145https://bit.ly/3oZoziI for Sinhala sentiment analysis was introduced by Jenarthanan et al. [138]. The dataset contains 318,308 Sinhala tweets annotated with emotions. The study by Dhananjaya et al. [301] uses a sentiment lexicon of a high‐resource language to fine-tune Pre‐trained multilingual language models (PMLMs) such as mBERT and XLM‐R on an intermediate task which in turn is then used on the sentiment classification task of Sinhala.
A sentiment-annotated dataset of Sinhala-English code mixed YouTube comments has been created by Uthpala and Thirukumaran [302]. However, given that even their paper is behind a paywall, there is no free and public way to access this dataset. The work by Gurgurov et al. [239] compared the performance of LLMs for the Sinhala sentiment analysis task. They report the highest results for the Masked Language Modeling objective for Sinhala in both of their configurations where they used a mixture of ConceptNet (CN) and/or Wikipedia (wiki) language adapters in conjunction with Task Adapters (TA). Liyanaarachchi et al. [303] used character-level embedding to classify sentiments of Sinhala news comments as positive or negative. They claim that the 1-of-m embedding outperforms Log-m embedding for this task. Bandaranayake and Usoof [304] used BERT, DistilBERT, RoBERTa, and XLM-R to analyse the sentiment of Sinhala news comments [257]. They report that XLMR-large model achieves the best result. Fernando and Ranathunga [305] in their study on improving cross-lingual representation of multilingual language models for low-resource languages using linguistic entity masking used Sinhala code mixed sentiment analysis as one of the evaluations. The work by Peiris and Sumanathilaka [306] analyses emotions in Sinhala songs both in the directions of lyrics and audio. In their analysis of lyrics, they use RoBERTa to represent the Sinhala words. Rizvi et al. [307] used XLM-RoBERTa to extract keywords and classify aspects in Sinhala Code-Mixed data. Then in their follow up work [308] they again used XLM-RoBERTa, this time for sentiment analysis and, on top of that used SHAP [309] and LIME [251] to make the model explainable. The short study on the sentiments of Sinhala comments by Mohamed et al. [310] reported best results overall to be from XLM-R and SinBERT, while BiLSTM and Logistic Regression showing reasonable results on certain tasks.
@@ 3.12 Hate Speech Detection
A machine learning approach to detect hate speech in Sinhala was proposed by De Silva [312]. A feature model and a data set146146146https://bit.ly/3FTNaMZ for Sinhala hate speech detection for youtube was proposed by De Saa and Ranathunga [131]Sandaruwan et al. [313] have attempted to identify abusive Sinhala comments in social media using text mining and machine learning techniques. A cyberbullying comment classification study for Sinhala was conducted by [314] where they used classical machine learning algorithms. The study by Hettiarachchi et al. [315] used classical machine learning methods to detect hate speech in Romanized Sinhala social media posts. While the basic idea is the same, they have avoided mentioning transliteration in their paper. The study by Samarasinghe et al. [316] proposed using CNNs for detecting hate speech in Sinhala text. Kariyawasam [317] proposed a machine learning approach for identifying toxic Sinhala language on social media. Guruge et al. [318] used an ensemble of Naïve Bayes, Support Vector Machine, XGBoost, MLP, and AdaBoost to detect hate speech in 49019 Tweets they collected from February of 2021 to April of 2021. Sandaruwan et al. [319] collected 3000 Sinhala comments corpus and conducted Multinomial Naïve Bayes to detect hate speech. Sheran [320] claims to have used machine learning to detect hate speech written in Sinhala or Singlish on social media. A deep learning approach for detecting hate speech in Sinhala tweets was explored by Munasinghe and Thayasivam [321]. The study by Shalinda and Munasinghe [322] utilized classical machine learning techniques such as Linear Support vector classifier, Random Forest Classification, SGD classifier, Logistic Regression, XGBoost classifier and multinomial Naive Bayes classifier on both Sinhala and Singlish (Romanized Sinhala) to identify hate words. They report that the SGD classifier over TF-IDF with uni-grams and bi-grams gives them the highest accuracy. Gamage et al. [323] conducted a comparative analysis on a number of embedding systems as well as classical frequency-based methods for Sinhala hate speech detection. The study by Fernando et al. [324] also claims to use machine learning and deep learning to detect hate speech in Sinhala. The study by Perera et al. [325] predicted Sinhala hate speech using user behaviour on twitter by applying ensemble Machine Learning to classify them by the five big personality traits. The follow-up study by Perera et al. [132] analyses Sinhala hate speech propagation on Twitter. They have also published a dataset147147147https://github.com/Isurie/Text-Classification-Module/tree/master/Dataset of 1600+ annotated Sinhala tweets. Rajapaksha et al. [326] used deep learning to identify trending periods of hate topics on Twitter in Sinhala. Arachchi et al. [327] used a web-based tool to translate between Sinhala and English to detect inappropriate word usage in Sinhala. The study by Ehelepola [328] analyzed Sinhala text from social media and e-commerce sites for hate speech. The work by Dikwatta et al. [329] used supervised algorithms to detect Sinhala hate speech in text from image posts on the Internet (memes). Later, they extended their work [330] by creating a Sinhala meme data set annotated with both the text and visual components. They then showed that their SVM-based method obtains 95% accuracy on this data set. Wickramaarachchi et al. [331] used LSTM on BART to compare the title and description of Sinhala youtube videos to their audio in order to determine if they contain hate speech. The study by Muthuthanthri and Smith [332] used an annotated dataset from Facebook with TF-IDF and fastText embedding features with SVM, CNN, RNN, LSTM as well as BERT and GPT2. They concluded that BERT obtains the best result for detecting hate speech in English-Sinhala code-mixed data. Chavinda and Thayasivam [174] have used Multilingual Large Language Models (MLLMs) with Dual Contrastive Learning (DCL) to detect Sinhala hate speech on Facebook148148148https://huggingface.co/datasets/krishan-CSE/Facebook_Sinhala_Hate_Speech and Twitter149149149https://huggingface.co/datasets/krishan-CSE/Twitter_Sinhala_Hate_Speech data sets. They report TwHIN-BERT-base [333] model to have obtained the best results. The work by Jahnavi and Chaturvedi [334] used BERT with CNN and LSTM to detect hate speech in Sinhala. Perera et al. [335] uses XLM-RoBERTa to classify Sinhala youtube video titles and comments to severity levels of hate. This, by definition is a multi-class classification. The work by Tanjim [336] analyses multiple low-resourced languages in the context of Cyberbullying detection. Ultimately, for the case of Sinhala as the target language, they recommend using Assamese, Bengali, Nepali, Urdu, Hindi, or Bodo as the source language for transfer learning this task.
@@
After the release of the SOLD dataset by Ranasinghe et al. [143] for the hate speech detection task in Sinhala, a number of subsequent works have been conducted using it. Fernando and Deng [337] used traditional machine learning algorithms to predict Sinhala hate speech. Ranasinghe and Zampieri [338] compared the results of pre-trained mT5 on Sinhala tweets from the SOLD dataset against results for German, Spanish, Hindi, Korean, and Marathi. The work by Bestgen [339] compared the result of classical machine learning algorithms on simple n-gram features for the hate speech detection task on the SOLD dataset for Sinhala against results for Assamese and Bengali. Dmonte et al. [340] used the SOLD data set to determine whether machine translation impacts the performance of hate speech detection. Haturusinghe et al. [341] introduces a series of models named Subasa that are tested on the SOLD dataset. Their trained PMLs150150150https://github.com/haturusinghe/subasa-plm and LLMs151151151https://github.com/haturusinghe/subasa-llm are separately publicly available. Ou6t of these models, they report that their Subasa-XLM-R model outperforms even the state-of-the-art commercial LLMs when evaluated on the SOLD data set. Ranasinghe et al. [342] used the SOLD dataset for the Hate Speech And Offensive Content (HASOC) identification task152152152https://hasocfire.github.io/hasoc/2023/task1.html they introduced. In completing the task, the study by Narayan et al. [343] compared the effectiveness of a number of pre-trained deep learning models for Sinhala hate speech detection. They showed that XLM Roberta Base achieves a macro-f1 of 83.49% over the LSTM baseline of 75.30%. For the same task, Ojo et al. [344] used mBERT to identify hate and offensive content in Sinhala. The work by Rostamkhani and Eetemadi [345] compared the zero-shot performance of ChatGPT on the SOLD dataset against a number of other Sinhala models. They conclude that Zero-shot ChatGPT perform better in Sinhala hate speech detection than models that were not fine-tuned on SOLD. Dmonte et al. [144] used the SOLD data set to verify the new MT-Offense153153153https://github.com/LanguageTechnologyLab/MT-Offense data set they created by using NMT to translate the English offensive language benchmark dataset, OLID [145]. They report the best F1 of 0.550 for the translation done using nllb-200-3.3B [137].
@@ 3.13 Fake News Detection
A dataset consisting of Sinhala documents drawn from Sri Lankan news websites was published by Jayawickrama et al. [129] along with the benchmark misinformation classification models. A hybrid approach to detect Sinhala fake news on Social media was proposed by Wijayarathna and Jayalal [346] where the text content of the post is checked against credible sources and the authenticity of the user account posting the relevant post is evaluated by a rule-based points allocation schema. Wijayarathna and Jayalal [347] collected a set of 120 fake news tweets and 250 non-fake news tweets which they then converted to vectors by taking the fasttext vector for words and averaging. The vector representations of the tweets were then compared to predict whether the news containing it is real or fake. The study by Udurawana et al. [348] proposes to use an accuracy score (obtained by analysing the news content) and a credibility score (obtained by using a scoring mechanism) to detect fake news in Sinhala text. They also incorporate a module that classifies on the basis of passive aggressiveness. The study by Adihetti and Jayalal [349] used autoencoders to detect Sinhala fake news from social media posts. The study by Wickramaarachchi et al. [331] detected fake content in Sinhala youtube videos by comparing their title and description to the audio using LSTM on BART. Atthanayake and Ruzaik [350] conducted a shallow comparison between the English Fakedit fake news data set [351] and the LIRNEasia Sinhala fake news data set [129].
@@ 3.14 Word Sense Disambiguation
There have been multiple attempts to do word sense disambiguation (WSD) [352, 353, 354, 355, 356] for Sinhala. For this, Arukgoda et al. [201] have proposed a system named Aruth based on the Lesk Algorithm[357]. An online tool154154154http://aruth.herokuapp.com/, an API155155155https://bit.ly/3sJEYbS of the algorithm, and code along with data on github156156156https://github.com/jseanm1/aruthSWSD are available. For the same task, Marasinghe et al. [358] have proposed a system based on probabilistic modeling. A dialogue act recognition system which utilizes simple classification algorithms has been proposed by Palihakkara et al. [359]. A word sense disambiguation tool named Sinsense was introduced by Subasingha [360]. They used cross-lingual sense disambiguation where English sense disambiguation was used to obtain Sinhala sense disambiguation. However, neither their tool nor their full research is publicly available.
@@ 3.15 Text Summarizing
A deterministic process flow for automatic Sinhala text summarizing was proposed by Welgama [361]. The study by Wimalasuriya [362], which has the same name as the above work by Welgama [361], uses the graph-based TextRank algorithm for automatic Sinhala text summarizing. The use case of Sinhala Text summarization for government gazettes was explored by Jayawardane [363]. The study by Rathnayake et al. [364] compared the results of extractive and abstractive summarization on Sinhala textbooks. The study by Jahan and Wijesekara [365] compared the abstractive summarizing methods of TF-IDF and Text-Rank for Sinhala using ROUGE as the evaluation score. They concluded TF-IDF to be the superior choice. Hasan et al. [154] introduced the XL-sum157157157https://github.com/csebuetnlp/xl-sum/tree/master dataset which includes 3,414 Sinhala documents and their summaries collected from Sinhala BBC website158158158https://www.bbc.com/sinhala. The M3LS159159159https://github.com/anubhav-jangra/M3LS multi-modal dataset introduced by Verma et al. [155] contains 10,148 Sinhala documents, relevant images, and their summaries collected from Sinhala BBC website160160160https://www.bbc.com/sinhala. Further, they claim that MT5 [366] obtains the best ROUGE-1 and ROUGE-L f scores for their Sinhala data set. Patabandi et al. [367] used BART and DistilBART to summarise Romanised Sinhala (Singlish) messages collected from WhatsApp group chats. The first Sinhala multi-document summarizing dataset, M2DS161161161https://huggingface.co/datasets/KushanH/m2ds, was created by Hewapathirana et al. [156]. The M2DS dataset consists of 23.5k Sinhala documents in 5.5k clusters with each cluster having a golden summary. Thier code and pre-trained models are also available on github162162162https://github.com/KushanMH/m2ds.